OpenStudy (anonymous):

Find f'(x) for f(x)=(x-4)/sqrt(2x-3)

OpenStudy (anonymous):

Do u know how to use the power rule? it allows you to calculate the derivative of f(x)

OpenStudy (anonymous):

i used chain rule for the bottom.

OpenStudy (anonymous):

so do you know the quotient rule. i assume its: \[\frac{x-4}{\sqrt{2x-3}}\]

OpenStudy (anonymous):

You can use the product rule as well by re-writing it as: \[(x-4)* \frac{1}{\sqrt{2x-3}} => (x-4)* \frac{1}{(2x-3)^{1/2}} => (x-4)(2x-3)^{-1/2}\]

OpenStudy (anonymous):

Btw, this is not differential equations :)

OpenStudy (anonymous):

i'd rather use quotient rule...

OpenStudy (anonymous):

oh sorry haha.

OpenStudy (anonymous):

so where are you stuck with the quotient rule

OpenStudy (anonymous):

anyhow, the quotient rule is easier for this :)

OpenStudy (anonymous):

i mean product rule** lol

OpenStudy (anonymous):

the combining them all together part...i used chain rule for the derivative of the bottom, and i got (-1/2(2x-3)^-3/2 + 2)

OpenStudy (anonymous):

times the top, so ((x-4)(-1/2(2x-3)^-3/2 + 2)-x+4)/2x-3

OpenStudy (anonymous):

That's after quotient rule and chain rule^

OpenStudy (anonymous):

so where are you stuck after it

OpenStudy (anonymous):

just the combining part. combining x-4 with the entire equation lol.

OpenStudy (anonymous):

can you use the drawing tool? or latex? I don't seem to decipher what you wrote..

OpenStudy (anonymous):

\[((x-4)(\frac{ 1 }{ 2 }(2x-3)^{\frac{ -3 }{ 2 }} +2)-x+4)/x-4\]

OpenStudy (anonymous):

It's the entire thing over x-4

OpenStudy (anonymous):

@Omniscience